OpenStudy (anonymous):

If tan x is -1, simplify tan(pi/3+x). Answer exact value and rationalize denominator.

OpenStudy (anonymous):

\[\tan(\frac{ π }{ 3 } + x) = \frac{ \tan(π/3) + \tan x }{ 1 - \tan(π/3)\tan x }\]\[= \frac{ \sqrt{3} + \tan x }{ 1 - \sqrt{3} \tan x}\]

OpenStudy (anonymous):

I'm sure you can rationalize it yourelf.

OpenStudy (anonymous):

And simplify too. Just plug.

OpenStudy (anonymous):

great, thank you

OpenStudy (anonymous):

would it be